### v1

# Models That Prove Their Own Correctness

Abstract
How can we trust the correctness of a learned model on a particular input of interest? Model accuracy is typically measured on average over a distribution of inputs, giving no guarantee for any fixed input. This paper proposes a theoretically-founded solution to this problem: to train Self-Proving models that prove the correctness of their output to a verification algorithm $V$ via an Interactive Proof.
Self-Proving models satisfy that, with high probability over a random input, the model generates a correct output and successfully proves its correctness to $V\!$. The soundness property of $V$ guarantees that, for every input, no model can convince $V$ of the correctness of an incorrect output. Thus, a Self-Proving model proves correctness of most of its outputs, while all incorrect outputs (of any model) are detected by $V$. We devise a generic method for learning Self-Proving models, and we prove convergence bounds under certain assumptions.
The theoretical framework and results are complemented by experiments on an arithmetic capability: computing the greatest common divisor (GCD) of two integers. Our learning method is used to train a Self-Proving transformer that computes the GCD and proves the correctness of its answer.
1 Introduction
Bob is studying for his algebra exam and stumbles upon a question $Q$ that he cannot solve. He queries a Large Language Model (LLM) for the answer, and it responds with a number: 42. Bob is aware of recent research showing that the LLM attains a 90% score on algebra benchmarks (cf. Frieder et al. 2023), but should he trust that the answer to his particular question $Q$ is indeed 42?
Bob could ask the LLM to explain its answer in natural language. Though he must proceed with caution, as the LLM might try to convince him of an incorrect answer (Turpin et al., 2023). Moreover, even if 42 is the correct answer, the LLM may fail to produce a convincing proof (Wang et al., 2023). If only the LLM could formally prove its answer, Bob would verify the proof and be convinced.
This paper initiates the study of Self-Proving models (Fig. 1) that prove the correctness of their answers via an Interactive Proof system (Goldwasser et al., 1985). Self-Proving models successfully convince a verification algorithm $V$ with worst-case soundness guarantees: for any question, with high probability over the interaction, $V$ will not be convinced of an incorrect answer. This is even when the prover with which $V$ is interacting has access to $V$’s specification, and far more computational power.
Figure 1: Self-Proving models. For input $x$, Self-Proving model $P_{\theta}$ generates an output $y$ and sends it to a Verification Algorithm $V$. Then, over $i\in[R]$ rounds, $V$ sends query $q_{i}$, and receives an answer $a_{i}$ from $P_{\theta}$. Finally, $V$ decides (“accept/reject”) whether it is convinced that $y$ is a correct output for $x$.
Our contributions are as follows.
  • We define Self-Proving models (Section 3).
  • We propose two methods for learning Self-Proving models in Section 4. The first, Transcript Learning (TL), relies on access to transcripts of accepting interactions and is the focus of this paper; we prove convergence bounds for TL under convexity and Lipschitzness assumptions. The second method, Reinforcement Learning from Verifier Feedback (RLVF), trains a model by emulating interaction with the verifier. We also present variants of these algorithms that use Annotations to improve learning in practice.
  • We empirically study the efficacy of TL and Annotated-TL (ATL) for training Self-Proving transformers that compute the Greatest Common Divisor (GCD) of two integers. Table 1 demonstrates the efficacy of our methods, with additional experiments in Section 5. Our results may be of independent interest for research on the arithmetic capabilities of transformers (e.g. Charton (2024); Lee et al. (2024)). Code, data and models will be made available upon publication.
Learning method Correctness Verifiability
GPT (baseline) $99.8\%$
GPT+TL $98.8\%$ $60.3\%$
GPT+ATL $98.6\%$ $96.7\%$
Table 1: Self-Proving transformers computing the GCD. We train a 6.7M parameter GPT to compute the GCD of two integers sampled log-uniformly from $[10^{4}]$. Vanilla GPT correctly generates the GCD for almost all inputs, but does not prove correctness to a simple verification algorithm. GPT trained with Transcript Learning (GPT+TL) proves its answer 60.3% of the time; training with Annotated Transcript Learning (GPT+ATL) increases this to 96.7%. See Section 5 for more details.
2 Related Work
This paper is situated at the intersection of machine learning (ML) theory and Interactive Proof systems (IPs). We briefly discuss recent relevant work from these literatures.
ML and IPs.
IPs have found numerous applications in ML towards a diverse set of goals. Anil et al. (2021) introduce Prover–Verifier Games, a game-theoretic framework for learned provers and verifiers. Wäldchen et al. (2024) cast the problem of model interpretability as a Prover–Verifier interaction between a learned feature selector and a learned feature classifier. Debate systems (Condon et al., 1995), a multiprover variant of IPs, were considered for aligning models with human values (Irving et al., 2018; Brown-Cohen et al., 2023). In such Debate systems, two competing models are each given an alleged answer $y\neq y^{\prime}$, and attempt to prove the correctness of their answer to a (human or learned) judge. Lastly, Murty et al. (2023) define Pseudointelligence: a model learner $L_{M}$ and an evaluator learner $L_{E}$ are each given samples from a ground-truth; $L_{M}$ learns a model of the ground-truth, while $L_{E}$ learns an evaluator of such models; the learned evaluator then attempts to distinguish between the learned model and the ground-truth in a Turing Test-like interaction.
All of these works consider learned verifiers, whereas our work focuses on training models that interact with a manually-defined verifier. More related in this regard is IP-PAC (Goldwasser et al., 2021), in which a learner proves that she learned a model that is Probably Approximately Correct (Valiant, 1984). We, however, consider models that prove their own correctness on a per-input basis, rather than learners that prove average-case correctness of a model.
Models that generate formal proofs.
Self-Proving models are verified by an algorithm with formal completeness and soundness guarantees (see Definition 3.2). In this sense, Self-Proving models generate a formal proof of the correctness of their output. Several works propose specialized models that generate formal proofs.
AlphaGeometry (Trinh et al., 2024) is capable of formally proving olympiad-level geometry problems; Gransden et al. (2015); Polu and Sutskever (2020); Yang et al. (2023) and others train models to produce proofs in Coq, Metamath and Lean (de Moura et al., 2015); FunSearch (Romera-Paredes et al., 2024) evolves LLM-generated programs by systematically evaluating their correctness. Indeed, all of these can be cast as Self-Proving models developed for specific proof systems. Meanwhile, this work defines and studies the class of such models in general. Several works (e.g. Welleck et al. 2022) consider models that generate natural language proofs or explanations, which are fundamentally different from formal proofs (or provers) verified by an algorithm.
Training on intermediate steps.
Chain-of-Though (CoT, Wei et al. 2022) refers to additional supervision on a model in the form of intermediate reasoning steps. CoT is known to improve model performance whether included in-context (Wei et al., 2022) or in the training phase itself (Yang et al., 2022). Transcript Learning (TL, Section 4.1) can be viewed as training the model on a Chain-of-Thought induced by the interaction of a verifier and an honest prover (Definition 3.2).
To complete the analogy, let us adopt the terminology of Uesato et al. (2022), who consider outcome supervision and process supervision. In our case, the outcome is the decision of the verifier, and the process is the interaction between the verifier and the model. Thus, Reinforcement Learning from Verifier Feedback (RLVF, Section 4.2) is outcome-supervised while TL is process-supervised. In a recent work, Lightman et al. (2024) find that process-supervised transformers outperform outcome-supervised ones on the MATH dataset (Hendrycks et al., 2021).
Transformers for arithmetic.
In Section 5 we train and evaluate Self-Proving transformers to generate the GCD of two integers and prove its correctness to a verifier. These experiments leverage a long line of work on neural models of arithmetic tasks originating with Siu and Roychowdhury (1992). Of particular relevance is the recent paper of Charton (2024), who trains transformers to generate the GCD—without a proof of correctness. We benefit from conclusions suggested in their work and start from a similar (scaled-down) experimental setup. Our main challenge (obtaining Self-Proving models) is overcome by introducing Annotated Transcript Learning (ATL).
We conduct ablation experiments to find two deciding factors in ATL. First, we study the effect of the amount of annotation given in the form of intermediate steps (Lee et al., 2024), which is related to (AR) length complexity (Malach, 2023). Second, we characterize ATL efficacy in terms of an algebraic property of the tokenization scheme (cf. Nogueira et al. 2021; Charton 2022, 2024).
3 Self-Proving models
We introduce and formally define our learning framework in which models prove the correctness of their output. We start with preliminaries from the learning theory and proof systems literatures in Section 3.1. We then introduce our main definition in Section 3.2.
3.1 Preliminaries
Let $\Sigma$ be a set of finite tokens and $\Sigma^{\ast}$ denote the set of finite sequences of such tokens. We consider sequence-to-sequence models $F_{\theta}\colon\Sigma^{\ast}\to\Sigma^{\ast}$, which are total functions that produce an output for each possible sequence. A model is parameterized by a real-valued, finite dimensional vector $\theta$. We consider models as randomized functions, meaning that $F_{\theta}(x)$ is a random variable over $\Sigma^{\ast}$, of which samples are denoted $y\sim F_{\theta}(x)$.
Before we can define models that prove their own correctness, we must first define correctness. Correctness is defined with respect to an input distribution $\mu$ over $\Sigma^{\ast}$, and a ground-truth $F^{\ast}$ that defines correct answers. For simplicity of presentation, we focus on the case that each input $x\in\Sigma^{\ast}$ has exactly one correct output $F^{\ast}(x)\in\Sigma^{\ast}$, and a zero-one loss function on outputs (the general case is left for future work). The fundamental goal of machine learning can be thought of as learning a model of the ground truth $F^{\ast}$. Formally,
Definition 3.1 (Correctness).
Let $\mu$ be a distribution of input sequences in $\Sigma^{\ast}$ and let $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$ be a fixed (deterministic) ground-truth function. The correctness of model $F_{\theta}$ with respect to $\mu$ is
$\mathrm{corr}(\theta)\coloneqq\mathrm{corr}(F_{\theta})\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim F_{\theta}(x)\end{subarray}}[y=F^{\ast}(x)].$
We say that model $F_{\theta}$ is $\alpha$-correct (on average) if $\mathrm{corr}(\theta)\geq\alpha$.
An interactive proof system (Goldwasser et al., 1985) is a protocol carried out between an efficient verifier and a computationally unbounded prover. The prover attempts to convince the verifier of the correctness of some assertion, while the verifier accepts only correct claims. The prover is powerful yet untrusted; in spite of this, the verifier must reject false claims with high probability.
In the context of this work, it is important to note that the verifier is manually-defined (as opposed to learned). Formally, the verifier is a probabilistic polynomial-time algorithm tailored to a particular ground-truth capability $F^{\ast}$. Informally, the verifier is the anchor of trust: think of the verifier as an efficient and simple algorithm, hosted in a trustworthy environment.
Given an input $x\in\Sigma^{\ast}$, the model $F_{\theta}$ “claims” that $y\sim F_{\theta}(x)$ is correct. We now define what it means to prove this claim. We will use $P_{\theta}$ to denote Self-Proving models, noting that they are formally the same object (a randomized mapping from $\Sigma^{\ast}$ to $\Sigma^{\ast}$) as vanilla models. We change the notation to emphasize that $P_{\theta}$ outputs $y\sim P_{\theta}(x)$ but can also be prompted by the verifier, unlike $F_{\theta}$ who is only expected to generate an output.
A Self-Proving model proves that $y\sim P_{\theta}(x)$ is correct to a verifier $V$ over the course of $R$ rounds of interaction (Figure 1). In each round $i\in[R]$, verifier $V$ queries $P_{\theta}$ on a sequence $q_{i}\in\Sigma^{\ast}$ to obtain an answer $a_{i}\in\Sigma^{\ast}$; once the interaction is over, $V$ accepts or rejects. For fixed $x,y\in\Sigma^{\ast}$, the decision of $V$ after interacting with $P_{\theta}$ is a random variable over $V$’s decision (accept/reject), determined by the randomness of $V$ and $P_{\theta}$. The decision random variable is denoted by $V^{P_{\theta}}(x,y)$.
We present a definition of Interactive Proofs restricted to our setting.
Definition 3.2.
Fix a soundness error $s\in(0,1)$, a finite set of tokens $\Sigma$ and a ground truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. A verifier $V$ (in an Interactive Proof) for $F^{\ast}$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover $P$ [^1] We intentionally write $P$ rather than $P_{\theta}$: Interactive Proofs are defined with respect to all possible provers, not just parameterized ones. It interacts with $P$ over $R$ rounds (see Figure 1) and outputs a decision $V^{P}(x,y)\in\{0,1\}$. Verifier $V$ satisfies the following two guarantees:
  • Completeness: There exists an honest prover $P^{\ast}$ such that, for all $x\in\Sigma^{\ast}$,
    $\Pr[V^{P^{\ast}}\!(x,F^{\ast}(x))\ \mathrm{accepts}]=1,$
    where the probability is over the randomness of $V$ [^2] WLOG, the honest prover is deterministic by fixing the optimal randomness of a randomized prover.
  • Soundness: For all $P$ and for all $x,y\in\Sigma^{\ast}$, if $y\neq F^{\ast}(x)$ then
    $\Pr[V^{P}(x,y)\ \mathrm{accepts}]\leq s.$
    where the probability is over the randomness of $V$ and $P$, and $s$ is the soundness error.
By definition, the soundness error $s$ of a verifier $V$ bounds the probability that it is mistakenly convinced of an incorrect output; in that sense, the smaller $s$, the “better” the verifier $V$. In our setting, we think of a manually-defined verifier $V$ who is formally proven (by a human) to have a small soundness error by analysis of $V$’s specification.
We remark that, as depicted in Figure 1, each of the model’s answers depends on all previous queries and answers in the interaction. This captures the setting stateful models, e.g. a session with a chatbot.
In anticipation of Self-Proving models (Section 3.2), let us observe the following. Completeness and soundness are worst-case guarantees, meaning that they hold for all possible inputs $x\in\Sigma^{\ast}$. In particular, completeness implies that for all $x\in\Sigma^{\ast}$, the honest prover $P^{\ast}$ convinces $V$ of the correctness of $F^{\ast}(x)$; in classical proof systems there is no guarantee that an “almost honest” prover can convince the verifier (cf. Paradise (2021)). Yet, if we are to learn a prover $P_{\theta}$, we cannot expect it to agree with $P^{\ast}$ perfectly, nor can we expect it to always output $F^{\ast}(x)$. Indeed, Self-Proving models will have a distributional guarantee with respect to inputs $x\sim\mu$.
3.2 Self-Proving models
We define the Verifiability of a model $P_{\theta}$ with respect to an input distribution $\mu$ and a verifier $V$. Intuitively, Verifiability captures the ability of the model to prove the correctness of its answer $y\sim P_{\theta}(x)$, when the input $x$ is sampled from $\mu$. We call models capable of proving their own correctness as Self-Proving models.
Definition 3.3 (Self-Proving model).
Fix a verifier $V$ for a ground-truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$ as in Definition 3.2, and a distribution $\mu$ over inputs $\Sigma^{\ast}$. The Verifiability of a model $P_{\theta}\colon\Sigma^{\ast}\to\Sigma^{\ast}$ is defined as
$\mathrm{ver}_{V,\mu}(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\end{subarray}}\left[V^{P_{\theta}}(x,y)\ \mathrm{accepts}\right].$ (1)
We say that model $P_{\theta}$ is $\beta$-Self-Proving with respect to $V$ and $\mu$ if $\mathrm{ver}_{V,\mu}(\theta)\geq\beta$.
Now, consider any input distribution $\mu$, ground-truth $F^{\ast}$, and a verifier $V$ for $F^{\ast}$ with soundness error $s$. By a union bound, if model $P_{\theta}$ is $\beta$-Verifiable, then it is $(\beta-s)$-correct. That is to say, Verifiability is formally a stronger guarantee than correctness when $V$ has small soundness error $s$.
The benefit of Verifiability over correctness is captured by the following scenario. Alice wishes to use a model $P_{\theta}$ to compute some functionality $F^{\ast}$ on an input $x_{0}$ in a high risk setting. Alice generates $y_{0}\sim P_{\theta}(x_{0})$. Should Alice trust that $y_{0}$ is correct? If Alice has a held-out set of labeled samples she can estimate $P_{\theta}$’s average correctness on $\mu$. Unfortunately, (average) correctness provides no guarantee regarding for the correctness of the particular $(x_{0},y_{0})$ that Alice has in hand. If, however, Alice has access to a verifier $V$ for which $P_{\theta}$ is Self-Proving, then she can trust the model on an input-by-input (rather than average-case) basis: Alice can execute $V$ on $(x_{0},y_{0})$ and black-box access to $P_{\theta}$. Soundness of $V$ guarantees that if $y_{0}$ is incorrect, then $V$ rejects with high probability, in which case Alice should either generate $P_{\theta}(x_{0})$ again—or find a better model.
4 Learning Self-Proving Autoregressive Models
With a sound verifier $V$ at hand, obtaining Self-Proving models with respect to $V$ holds great promise: a user that prompts the model with input $x$ does not need to take it on good faith that $P_{\theta}(x)$ is correct; she may simply verify this herself by executing the verification protocol. How, then, can we learn models that are not just approximately-correct, but Self-Proving as well?
The challenge is to align the model with a verifier. We assume that the learner has access to input samples $x\sim\mu$ and correct outputs $F^{\ast}(x)$, as well as the verifier specification (code). Additionally, the learner can emulate the verifier, as the latter is required to be computationally efficient. [^3] We refer the reader to classical literature on Interactive Proof systems for formal definitions of computational efficiency (e.g. Goldreich (2008)).
Our focus is on autoregressive sequence-to-sequence (Self-Proving) models $P_{\theta}$ (Elman, 1990). Such models generate their output by recursively prompting a randomized sampling from a base distribution $p_{\theta}$ over tokens $\Sigma$. For an input $z\in\Sigma^{\ast}$, the output $w\sim P_{\theta}(z)$ is generated as follows:
  • Sample $w_{1}\sim p_{\theta}(z)$.
  • Let $j=1$. While $w_{j}$ is not the end-of-sequence token $\mathtt{EOS}\in\Sigma^{\ast}$:
    • Sample $w_{j+1}\sim p_{\theta}(zw_{1}\cdots w_{j})$.
  • Output $w=w_{1}w_{2}\cdots w_{j}$.
For any $z\in\Sigma^{\ast}$, it is useful to consider the vector of log-proabilities over $\Sigma$, denoted by $\log p_{\theta}(z)\in\mathbb{R}^{|\Sigma|}$. We assume that each coordinate in this vector is differentiable with respect to $\theta$.
Our general approach is inspired by Reinforcement Learning from Human Feedback (Christiano et al., 2017), a method for aligning models with human preferences, which has recently been used to align sequence-to-sequence models (Ouyang et al., 2022). However, there are two important differences between humans and algorithmic verifiers: (1) Verifiers are efficient algorithms which may be emulated by the learner. This is unlike humans, whose preferences are costly to obtain. On the other hand, (2) verifiers make a single-bit decision at the end of an interaction, but cannot guide the prover (model) in intermediate rounds. In RL terms, this is known as the exploration problem for sparse reward signals (e.g., Ladosz et al. 2022).
Section 4.1 introduces Transcript Learning (TL), a learning algorithm that overcomes the exploration problem mentioned in the second point under the assumption that the learner has access to transcripts of interactions in which the verifier accepts. We prove convergence bounds for TL (Section A.1) and analyze it experimentally (Section 5).
Access to accepting transcripts is a reasonable assumption, for example, when there is an efficient honest prover that can generate such transcripts (Goldwasser et al., 2015). When there is no access to accepting transcripts, we propose Reinforcement Learning from Verifier Feedback (Section 4.2).
4.1 Transcript Learning
We present an algorithm for learning Self-Proving models which uses access to a distribution of accepting transcripts. This is a reasonable assumption to make when the honest prover $P^{\ast}$ (see Definition 3.2) is efficient, as is the case in Doubly-Efficient Interactive Proof systems as defined by Goldwasser et al. (2015) and developed in other theoretical (e.g. Goldreich and Rothblum 2018) and applied (e.g. Zhang et al. 2021) works. In this case, an honest prover $P^{\ast}$ can be run by the learner during training to collect accepting transcripts without incurring heavy computational cost.
The intuition behind Transcript Learning is that the interaction of the verifier and prover can be viewed as a sequence itself, which is called the transcript $\pi\in\Sigma^{\ast}$. The idea is to learn a model not just of $x\mapsto y^{\ast}$ for a correct output $y^{\ast}$, but of $x\mapsto y^{\ast}\pi^{\ast}$, where $\pi^{\ast}$ is a transcript of an interaction in which the verifier accepted.
In more detail, Transcript Learning requires access to an (honest) transcript generator $\mathcal{T}^{\ast}$. Given an input $x$, the generator $\mathcal{T}^{\ast}(x)$ samples a sequence $P^{\ast}\!(x)\pi^{\ast}\in\Sigma^{\ast}$ such that $\pi^{\ast}$ is an accepted transcript. The generator is autoregressive, meaning that for any prefix of an accepted transcript $\pi^{\ast}_{\leq t}\in\Sigma^{t}$, the learner has access to the distribution over next tokens $\mathcal{T}^{\ast}(\pi_{\leq t})\in\Sigma$ [^4] Formally, if the generator is prompted with any string that cannot be completed to an accepted transcript, it outputs a dummy symbol $\bot\in\Sigma$.
Transcript Learning (TL) trains a Self-Provable model by autoregressively optimizing towards generating accepting transcripts. It is described in Algorithm 1. At a very high level, it works by repeatedly sampling $x\sim\mu$ and transcript $y^{\ast}\pi^{\ast}\sim\mathcal{T}^{\ast}(x)$, and updating the logits $\log p_{\theta}$ towards agreeing with $y^{\ast}\pi^{\ast}$ via Gradient Ascent. We prove that, under certain conditions, it is expected to output a Self-Provable model. We fully describe the conditions in Section A.1; we provide an informal statement (with an undefined “expected agreement” quantity) next.
Theorem 4.1 (Theorem A.5, informal).
Fix an input distribution $\mu$, a verifier $V$ and an autoregressive model family $\{P_{\theta}\}_{\theta}$. Fix a transcript generator $\mathcal{T}^{\ast}$ such that expected agreement $\mathcal{T}^{\ast}$ is convex in $\theta$. For any $\varepsilon>0$ such that there exists $\theta^{\ast}$ with at least $\geq 1-\varepsilon/2$ expected agreement with $\mathcal{T}^{\ast}$, let $B$ be the minimal norm of such $\theta^{\ast}$. Let $\rho>0$ such that for all $\theta$ with $||\theta||<B$, the logits $\log p_{\theta}$ are $\rho$-Lipschitz in $\theta$. Denote by $\bar{\theta}$ the output of TL running for number of iterations
$N=R^{2}\cdot(L_{a}+1)^{2}\cdot\frac{4\rho^{2}\cdot B^{2}}{\epsilon^{2}}$ (2)
and learning rate $\lambda=\varepsilon/2R^{2}L_{a}^{2}\rho$. Then the expected Verifiability of $\bar{\theta}$ is at least $1-\varepsilon$.
The proof works by reduction to Stochastic Gradient Descent (SGD). The main challenge is to prove that the learner can use its only available tools—sampling honest transcripts, emulating the verifier, and differentiating the logits—to estimate the gradient of (a lower bound on) the Verifiability of $P_{\theta}$.
Looking at Equation 2, we see that the sample complexity of TL grows like that of SGD (e.g. Shalev-Shwartz and Ben-David 2014), multiplied by the number of rounds and length of answers in the proof system. Minimizing these quantities (known collectively as the communication complexity) has been an overarching goal in the study of proof systems (e.g. Goldreich and Håstad (1998); Goldreich et al. (2002); Reingold et al. (2021). Theorem 4.1 formally shows the benefit of communication-efficient proof systems in the context of Self-Proving models.
4.2 Reinforcement Learning from Verifier Feedback (RLVF)
As mentioned in Section 4.1, Transcript Learning uses access to an honest transcript generator to estimate gradients of (a lower bound on) the Verifiability of a model $P_{\theta}$.
Reinforcement Learning from Verifier Feedback (RLVF, Algorithm 2) estimates this gradient without access to a transcript generator. RLVF can be viewed as a modification of TL in which the learner emulates the interaction of the verifier with its own model $P_{\theta}$. Rather than directly sampling from the generator as in TL, it collects accepting transcripts by rejection sampling on emulated transcripts.
This rejection sampling means that RLVF requires its initial model $P_{\theta_{0}}$ to have Verifiability bounded away from 0, so that accepting transcripts are sampled with sufficient probability. Fortunately, such a Self-Proving base model can be learned using TL. This gives a learning paradigm in which a somewhat-Self-Proving base model is first learned with TL (with Verifiability $\delta>0$), and then “amplified” to a fully Self-Proving model using RLVF (cf. Nair et al. 2018).
We prove that RLVF learner can estimate the Verifiability gradient of $P_{\theta}$ using emulation alone in Lemma A.6. From a broader perspective, RLVF can be derived by viewing Self-Proving as a reinforcement learning problem in which the agent (prover) is rewarded when the verifier accepts. Indeed, RLVF is the Policy Gradient method (Sutton et al., 1999) for a verifier-induced reward. Convergence bounds for Policy Gradient methods are a challenging and active area of research (e.g. Agarwal et al. 2021), and so we leave the full analysis to future work.
4.3 Learning from annotated transcripts
To minimize the length of messages exchanged in an Interactive Proof system, the honest prover is designed to send the shortest possible message to the verifier, containing only essential information.
However, when training Self-Proving model, it may be useful for it to first generate an “annotated” answer $\widetilde{a}$ which is then trimmed down to the actual answer $a$ to be sent to the verifier. We formally adapt the framework from Sections 3 and 4 to this setting in Appendix B, where we present Annotated Transcripts. This can be viewed as adding Chain-of-Thought (Wei et al., 2022) to the model. The Transcript Learning algorithm naturally extends to annotated transcripts as well.
5 Experimental Results
We describe our experimental setup, and present ablation studies that shed additional light on the effect of annotation and representation on Verifiability.
5.1 Setup: Training transformers to predict the GCD of two integers
Charton (2024) empirically studies the power and limitations of learning GCDs with transformers. We follow their setup and two conclusions on settings that make for faster learning: Training from the log-uniform distribution, and choosing a base of representation with many prime factors.
[…]
We choose the input distribution $\mu$ to be the log-uniform distribution on $[10^{4}]$, and train the transformer on sequences of the form $\mathbf{(x_{1}x_{2}y)}$, where $x_{1},x_{2}\sim\mu$ and $y=GCD(x_{1},x_{2})$. This is a scaling-down of Charton (2024), to allow single GPU training of Self-Proving transformers. In all of our experiments, we use a GPT model (Vaswani et al., 2017) with 6.3M parameters trained on a dataset of 1024K samples in batches of 1024. Full details are deferred to Appendix D.
Proving correctness of GCD.
Following Charton (2024) as a baseline, we find that transformers can correctly compute the GCD with over $99\%$ probability over $(x_{1},x_{2})\sim\mu$. To what extent can they prove their answer? To answer this question, we first devise a natural proof system based on Bézout’s theorem. Its specification and formal guarantees are deferred to Appendix C. We denote its verification algorithm by $V$, and highlight some important features of the experimental setup:
  • The proof system consists of one round ($R=1$). The verifier makes no query, and simply receives a proof $\mathbf{\pi}$ from the prover.
  • Completeness: For any $x_{1},x_{2},y\in[10^{4}]$ such that $y=GCD(x_{1},x_{2})$, there exists a proof $\mathbf{\pi}$ such that $V(\mathbf{x_{1}x_{2}y\pi})$ accepts. As detailed in Appendix C, the proof $\pi$ consists of a pair of integers who are Bézout coefficients for $x_{1},x_{2}$.
  • Soundness: If $y\neq GCD(x_{1},x_{2})$, then $V(\mathbf{x_{1}x_{2}y\pi})$ rejects for any alleged proof $\mathbf{\pi}\in\Sigma^{\ast}$.
To measure Verifiability, we train a Self-Proving transformer using Transcript Learning on sequences $\mathbf{(x_{1}x_{2}y\pi)}$ and estimate for how many inputs $x_{1},x_{2}\sim\mu$ does the model generate both the correct GCD $\mathbf{y}$ and a valid proof $\mathbf{\pi}$. We test on 1000 pairs of integers $x_{1}^{\prime},x_{2}^{\prime}\sim\mu$ held-out of the training set, prompting the model with $\mathbf{(x_{1}^{\prime}x_{2}^{\prime})}$ to obtain $\mathbf{(y^{\prime}\pi^{\prime})}$, and testing whether $V(\mathbf{x_{1}^{\prime}x_{2}^{\prime}y^{\prime}\pi^{\prime}})$ accepts.
[…]
5.2 Models generalize beyond annotations
Figure 2: Verifiability with increasing amounts of annotation. $T$ is the number of steps added in Annotated Transcript Learning. Dashed lines bound the Verifiability of models that can only prove for integers up to a certain number of steps. Off chart to the left are bounds for depths 3 (47%), 4 (63%), and 5 (75%). Each $T$ was run with three seeds, with mean $\pm$ standard error depicted. See Appendix E for additional figures.
The proof $\mathbf{\pi}$ is annotated by including intermediate steps in its computation. Details are deferred to Appendix C; roughly speaking, we observe that the proof $\mathbf{\pi}$ for input $\mathbf{(a,b)}$ is obtained as the last element in a sequence $\mathbf{a,b,\pi_{1},\pi_{2},\dots}$ computed by the Euclidean algorithm. We annotate the proof $\mathbf{\pi}$ by prepending to it the sequence of Euclidean steps $\mathbf{(\pi_{1},\dots,\pi_{T})}$ up to some fixed cutoff $T$.
Figure 2 shows how $T$ affects the Verifiability of the learned model. As suggested by Lee et al. (2024), training the model on more intermediate steps results in better performance; in our case, increasing the number of intermediate steps $T$ yields better Self-Proving models. One might suspect that models only learn to execute the Euclidean algorithm in-context. To rule out this hypothesis, we derive an upper bound on the possible efficacy of such limited models. This bound is based on the Euclidean depth of integers $(x_{1},x_{2})$, which we define as the number of intermediate steps that the Euclidean algorithm makes before terminating on input $(x_{1},x_{2})$. Indeed, a model that only learns the to compute (in-context) the simple arithmetic of the Euclidean algorithm would only be able to prove the correctness of inputs $(x_{1},x_{2})$ whose depth does not exceed the annotation cutoff $T$.
Figure 2 tells a different story: For each cutoff $T$, we estimate the probability that integers $x_{1},x_{2}\sim\mu$ have Euclidean depth at most $T$ on $10^{5}$ sampled pairs. Larger annotation cutoff $T$ increases Verifiability, but all models exceed their corresponding Euclidean depth bound.
5.3 Base of representation
Figure 3: The number of prime divisors of a base $\omega(B)$ determines Verifiability. For each $o\in[4]$, we sampled 17 bases $B\in\{2,\dots,1386\}$ such that $\omega(B)=o$. A Self-Proving transformer was trained via Transcript Learning for twenty epochs on an identical dataset of 1024K samples encoded in base $B$. For each $\omega(B)$ we depict the mean $\pm$ standard error.
As mentioned previously, Charton (2024) concludes that, for a given base of representation $B$, transformers correctly compute the GCD of integers $x_{1},x_{2}$ that are products of primes dividing $B$. Simply put, choosing a base $B$ with many different prime factors yields models with better correctness (accuracy), which suggests why base $B=210=2\cdot 3\cdot 5\cdot 7$ yielded the best results.
To test whether the factorization of $B$ has a similar effect on Verifiability as well, we train transformers on 68 bases varying the number of prime divisors $\omega(B)$ from $\omega(B)=1$ (i.e., $B$ is a prime power) to $\omega(B)=4$. Figure 3 shows that $\omega(B)$ correlates not just with correctness (Charton, 2024), but also with Verifiability. Although the finding is statistically significant (no overlapping error margins), the overall difference is by a few percentage points; we attribute this to the smaller (10%) number of samples on which models were trained, relative to our other experiments.
6 Conclusions
Trust between a learned model and its user is fundamental. In recent decades, Interactive Proofs (Goldwasser et al., 1985) have emerged as a general theory of trust established via verification algorithms. This work demonstrates that models can learn to formally prove their answers in an Interactive Proof system. We call models that possess this capability Self-Proving.
The definition of Self-Proving models forms a bridge between the rich theory of Interactive Proofs and the contemporary topic of Trustworthy ML. Interactive Proofs offer formal worst-case soundness guarantees; thus, users of Self-Proving models can be confident when their models generate correct answers—and detect incorrect answers with high probability.
We demonstrate the theoretical viability of our definition with two generic learning algorithms: Transcript Learning (TL) and Reinforcement Learning from Verifier Feedback (RLVF). The analyses of these algorithms is informed by techniques from theories of learning, RL, and computational complexity. This work can be extended in several directions: finding conditions for the convergence of RLVF, improving sample complexity bounds for TL, or designing altogether different learning algorithms (for example, by taking advantage of properties of the verifier).
To better understand the training dynamics of (Annotated) TL, we train Self-Proving transformers for the Greatest Common Divisor (GCD) problem. We train a small (6.3M parameter) transformer that learns to generate correct answers and proofs with high accuracy. Facing forward, we note that Interactive Proofs exist for capabilities far more complex than the GCD (Shamir, 1992); scaling up our experiments is the next step towards bringing Self-Proving models from theory to practice.
Acknowledgments and Disclosure of Funding
We are grateful to Micah Carroll for helpful comments. This research was supported by DARPA-TA1 under grant no. HR001119S0076, and by the Simons Collaboration on the Theory of Algorithmic Fairness.
Appendix A Theoretical analyses for Section 4
In this section we provide a formal description and analysis of Transcript Learning (TL, Section 4.1) and Reinforcement Learning from Verifier Feedback (RLVF, Section 4.2). In Section A.1 we prove a convergence theorem for TL under convexity and Lipschitzness assumptions. Obtaining an analogous result for RLVF is more challenging; in lieu of a full analysis, we provide a lemma showing that the gradients estimated in the algorithm approximate the Verifiability of the model.
Specification of the learning model.
We must first fully specify the theoretical framework in which our results reside. Continuing from Section 3, we define a learner as an algorithm $\Lambda$ with access to a family of autoregressive models $\{P_{\theta}\}_{\theta}$ and samples from the input distribution $x\sim\mu$. In our setting of Self-Proving models (and in consistence with the Interactive Proof literature), we give the learner the full specification of the verifier $V$. More formally,
Definition A.1 (Self-Proving model learner).
A (Self-Proving model) learner is a probabilistic oracle Turing Machine $\Lambda$ with the following access:
  • A family of autoregressive models $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$ where $d\in\mathbb{N}$ is the number of parameters in the family. Recall (Section 4) that for each $\theta$ and $z\in\Sigma^{\ast}$, the random variable $P_{\theta}(z)$ is determined by the logits $\log p_{\theta}(z)\in\mathbb{R}^{|\Sigma|}$. For any $z\in\Sigma^{\ast}$ and $\sigma\in\Sigma$, the learner $\Lambda$ can compute the gradient of the $\sigma^{\text{th}}$ logit, that is, $\nabla_{\theta}\log\Pr_{\sigma^{\prime}\sim p_{\theta}(z)}[\sigma=\sigma^{\prime}]$.
  • Sample access to a the input distribution $\mu$. That is, $\Lambda$ can sample $x\sim\mu$.
  • The full specification of the verifier $V$, i.e., the ability to emulate the verification algorithm $V$. More specifically, $\Lambda$ is able to compute $V$’s decision after any given interaction; that is, given input $x$, output $y$, and a sequence of queries and answers $(q_{i},a_{i})_{i=1}^{R}$, the learner $\Lambda$ can compute the decision of $V$ after this interaction.
We remark that analysis of Transcript Learning will require a slight strengthening of the final item above. This is discussed in Section A.1.
Throughout this section, we will refer to the transcript of an interaction between a verifier and a prover (see Figure 1). We will denote $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$, and for any index $s\in|\pi|$ we will write $\pi_{<s}\in\Sigma^{s-1}$ to denote the $s$-long prefix of $\pi$. Throughout this section, we will use $\pi\in\Sigma^{\ast}$ to denote the transcript of an interaction between a verifier and a prover.
A.1 Transcript Learning
Recall that Transcript Learning requires access to an honest transcript generator. Before we can formally define this object, it will be useful to define a query generator for a verifier $V$.
Definition A.2 (Query generator).
Fix a verifier $V$ in a proof system with $R\in\mathbb{N}$ rounds, where the verifier issues queries of length $L_{q}=|q_{i}|$ and the prover (model) responses with answers of length $L_{a}=|a_{i}|$. [^6] We can assume that queries (resp. answers) all have the same length by padding shorter ones. The query generator $V_{q}$ corresponding to $V$ takes as input a partial interaction and samples from the distribution over next queries by $V$. Formally, for any $r\leq R$, given input $x$, output $y$, and partial interaction $(q_{i},a_{i})_{i=1}^{r}$, $V_{q}(x,y,q_{1},a_{1},\dots,q_{r},a_{r})$ is a random variable over $\Sigma^{L_{q}}$. [^7] For completeness’ sake, we can say that when prompted with any sequence $z$ that does not encode an interaction, $V_{q}(z)$ is fully supported on a dummy token $\bot\in\Sigma$.
We assume that access to the verifier specification (Definition A.1) includes access to the query generator. After all, the verifier—who is assumed to be efficient—sampled from $V_{q}$ during the interaction. Moreover, we will assume that for any partial interaction and any sequence $q^{\prime}$, the learner is able to compute the probability that $q^{\prime}$ was the next query. In other words, we assume the learner can compute the probability density function of $V_{q}$.
A transcript generator is a random variable over transcripts that faithfully represents the interaction of the verifier with some prover for a given input. An honest transcript generator is one who is fully supported on transcripts accepted by the verifier.
Definition A.3 (Honest transcript generator).
Fix a verifier $V$ in a proof system of $R\in\mathbb{N}$ rounds. A transcript generator $\mathcal{T}_{V}$ for $V$ is a randomized mapping from inputs $x\in\Sigma^{\ast}$ to transcripts $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})\in\Sigma^{\ast}$. For any input $x$, $\mathcal{T}_{V}(x)$ satisfies that for each $r\leq R$, the marginal of $\mathcal{T}_{V}(x)$ on the $r^{\text{th}}$ query ($q_{r}$) agrees with the corresponding marginal of the query generator $(V_{q})_{r}$.
A transcript generator $\mathcal{T}^{\ast}_{V}\coloneqq\mathcal{T}_{V}$ is honest if it is fully supported on transcripts $\pi^{\ast}$ for which the verifier accepts. [^8] WLOG we can assume that the prover sends her final answer $a_{R}$, the verifier’s decision is deterministic.
Notice that for any verifier $V$, there is a one-to-one correspondence between transcript generators and (possibly randomized) provers. We intentionally chose not to specify a prover in Definition A.3 to emphasize that transcripts can be “collected” independently of the honest prover (see completeness in Definition 3.2). As long as the generator is fully supported on honest transcripts, it can be used for Transcript Learning, described next (TL, Algorithm 1).
Hyperparameters: Learning rate $\lambda\in(0,1)$ and number of samples $N\in\mathbb{N}$.
Input: An autoregressive model family $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$, verifier specification (code) $V$, and sample access to an input distribution $\mu$ and an accepting transcript generator $\mathcal{T}_{V}^{\ast}(\cdot)$.
Output: A vector of parameters $\bar{\theta}\in\mathbb{R}^{d}$.
1
1:Initialize $\theta_{0}\coloneqq\vec{0}$.
22:for $i=0,\dots,N-1$ do
3:Sample $x\sim\mu$ and $\pi^{*}=(y,q_{1},a_{1},\dots,q_{R},a_{R})\sim\mathcal{T}_{V}^{\ast}(x)$. Denote $a_{0}\coloneqq y$.
3      4:foreach r = 0,…, R do
4            5:Let $S(r)$ denote the indices of the $r^{\text{th}}$ answer $a_{r}$ in $\pi^{\ast}$. for $s\in S(r)$ do
5                  6:Compute # Forwards and backwards pass
$\alpha_{s}(\theta_{i})$ $\coloneqq\Pr_{\sigma\sim p_{\theta_{i}}(x\pi_{<s})}[\sigma=\pi_{s}]$
$\vec{d}_{s}(\theta_{i})$ $\coloneqq\nabla_{\theta}\alpha_{s}(\theta_{i})=\nabla_{\theta}\log\Pr_{\sigma\sim p_{\theta_{i}}(x\pi_{<s})}[\sigma=\pi_{s}].$
7:If $r\geq 1$, let $q_{r}$ denote the $r^{\text{th}}$ query $q_{r}$ in $\pi^{\ast}$, and let $t$ denote its first index. That is, $\pi_{<t}^{\ast}=(y,q_{1},a_{1},\dots,q_{t-1},a_{t-1})$. Compute # Emulate the verifier
$\beta_{r}(\theta_{i})\coloneqq\Pr_{q^{\prime}\sim V_{q}(x\pi^{\ast}_{<t})}[q^{\prime}=q].$
6            8:Update
$\theta_{i+1}\coloneqq\theta_{i}+\lambda\cdot\alpha_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}b_{r}(\theta_{i})a_{s}(\theta_{i})\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})$
9:Output $\bar{\theta}\coloneqq\frac{1}{N}\sum_{i\in[N]}\theta_{i}$.
Algorithm 1 Transcript Learning (TL)
Convergence of TL is proven by a reduction to Stochastic Gradient Descent (SGD). Essentially, we are tasked with proving that TL estimates the gradient of the Verifiability of its model $P_{\theta}$. More precisely, TL estimates the gradient of a function that bounds the Verifiability from below. Maximizing this function therefore maximizes the Verifiability.
The lower-bounding function is the agreement of the transcript generator induced by $P_{\theta}$ with the provided honest transcript generator $\mathcal{T}^{\ast}_{V}$. More formally, we let $\mathcal{T}^{\theta}_{V}$ denote the transcript generator induced by the model $P_{\theta}$: for each $x$, $\mathcal{T}^{\theta}_{V}(x)$ is simply the distribution over transcripts of interactions between $V$ and $P_{\theta}$ on input $x$. We first prove that this function is differentiable.
Lemma A.4.
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. Fix an honest transcript generator $\mathcal{T}_{V}^{\ast}$. For any model $P_{\theta}$, it holds that
$\nabla_{\theta}\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right]=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\alpha_{0}(\theta)\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta)\cdot\alpha_{s}(\theta)\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)\right]$
where $S(r)$, $\beta_{r}(\theta)$, $\alpha_{s}(\theta)$ and $\vec{d}_{s}(\theta)$ are as defined in Algorithm 1.
Note that Lemma A.4 is true for any model $P_{\theta}$. Moreover, the random vector over which the expectation is taken (in the right hand side) is precisely the direction of the update performed in Algorithm 1. We now prove Lemma A.4, from which we derive Theorem 4.1.
Proof of Lemma A.4.
Throughout this proof, expectations and probabilities will be over the same distributions as in the lemma statement. First,
$\Pr_{x,\pi^{*},\pi}\left[\pi=\pi^{*}\right]=\operatorname*{\mathbb{E}}_{x,\pi^{*}}\Pr_{\pi}[\pi=\pi^{*}],$
and so, by the linearity of the gradient,
$\nabla_{\theta}\Pr_{x,\pi^{*},\pi}\left[\pi=\pi^{*}\right]=\operatorname*{\mathbb{E}}_{x,\pi^{*}}\nabla_{\theta}\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right).$ (3)
The probability that the output of $V$ and $P_{\theta}$ on input $x$ is equal to a given transcript is (by the law of total probability) the product of probabilities that each of the tokens of the transcript is equal to the corresponding token of the given transcript, both tokens generated by $V$’s queries and by $P_{\theta}$’s answers, when conditioning on the prefix of the transcript.
Formally, consider any fixed $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$ and denote the random variable $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$. For any $r\in[R]$ denote the random variables $V_{q}^{<r}\coloneqq V_{q}(y,q_{1},a_{1},\dots,q_{r-1},a_{r-1})$ and $\mathcal{T}^{\theta,<r}_{V}\coloneqq\mathcal{T}^{\theta}_{V}(yq_{1}a_{1}\cdots a_{r-1}q_{r})$. Then,
$\Pr_{\pi}\left[\pi=\pi^{*}\right]$ $\coloneqq\Pr_{\pi}[(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})=(y,q_{1},a_{1},\dots,q_{R},a_{R})]$ (4)
$=\Pr_{y\sim P_{\theta}(x)}[y=y^{\ast}]\cdot\prod_{\begin{subarray}{c}r\in[R]\end{subarray}}\Pr_{q\sim V_{q}^{<r}}[q=q_{r}^{\ast}]\cdot\Pr_{a\sim\mathcal{T}^{\theta,<r}_{V}}[a=a_{r}^{\ast}]$
$=\Pr_{y\sim P_{\theta}(x)}[y=y^{\ast}]\cdot\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\Pr_{q\sim V_{q}^{<r}}[q=q_{r}^{\ast}]\cdot\Pr_{\sigma\sim p_{\theta}(\pi^{\ast}_{<s})}[\sigma=\pi^{\ast}_{s}]$ (5)
$=\alpha_{0}(\theta)\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta)\cdot\alpha_{s}(\theta)\right),$ (6)
where Equation 4 uses independence of the verifier and models’ randomness, Equation 5 uses the autoregressive property of $P_{\theta}$ (Definition A.1), and Equation 6 is by definition of $\alpha_{s}$ and $\beta_{r}$.
Next, a basic calculus identity gives
$\nabla_{\theta}\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right)=\Pr_{\pi}\left[\pi=\pi^{*}\right]\cdot\nabla_{\theta}\log\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right).$ (7)
Let us focus on the last term. By Eq. 6,
$\nabla_{\theta}\log\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right)$ $=\nabla_{\theta}\log\alpha_{0}(\theta)\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta)\cdot\alpha_{s}(\theta)\right)$
$=\nabla\log_{\theta}\alpha_{0}(\theta)+\sum_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\nabla_{\theta}\log\beta_{r}(\theta)+\nabla_{\theta}\log_{\theta}\alpha_{s}(\theta)$
$=\nabla\log_{\theta}\alpha_{0}(\theta)+\sum_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\nabla_{\theta}\log_{\theta}\alpha_{s}(\theta)$ (8)
$=\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\nabla_{\theta}\log_{\theta}\alpha_{s}(\theta)=\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)$ (9)
where Equation 8 is because $\log\beta_{r}(\theta)\coloneqq\log\Pr_{q^{\prime}\sim V_{q}(x\pi^{\ast}_{<t})}[q^{\prime}=q]$ is a constant and therefore has a gradient of zeros, and Equation 9 is by definition of $\vec{d}_{s}(\theta)$.
Combining Equations 6, 7 and 9 concludes the proof. ∎
We are now ready to prove Theorem 4.1. We restate it below in full formality.
Theorem A.5.
[Theorem 4.1, formal] Fix a verifier $V$, an input distribution $\mu$, and an autoregressive model family $\{P_{\theta}\}_{\{\theta\in\mathbb{R}^{d}\}}$, and a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$. Fix an honest transcript generator $\mathcal{T}^{\ast}_{V}$ such that the expected agreement
$\mathrm{agree}_{\mathcal{T}^{\ast}_{V}}(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right]$
is convex in $\theta$. For any $\varepsilon>0$, we define
$B_{\varepsilon}$ $\coloneqq\min\left\{||\theta^{\ast}||:\mathrm{agree}_{\mathcal{T}^{\ast}_{V}}(\theta^{\ast})\geq 1-\varepsilon/2\right\}$
$\rho_{\varepsilon}$ $\coloneqq\max\left\{||\nabla_{\theta}\log p_{\theta}(z)||:z\in\Sigma^{\ast},||\theta||<B_{\varepsilon}\right\}$
For any $\varepsilon>0$ such that $B_{\varepsilon}$ and $\rho_{\varepsilon}$ are both finite, we denote by $\bar{\theta}$ the output of TL (Algorithm 1) running for number of iterations
$N=\frac{4R^{2}\cdot(L_{a}+1)^{2}\cdot\rho^{2}_{\varepsilon}\cdot B^{2}_{\varepsilon}}{\epsilon^{2}}$ (10)
and learning rate $\lambda=\varepsilon/2R^{2}L_{a}^{2}\rho$. Then the expected Verifiability (over the randomness of the samples collected by TL) is at least $1-\varepsilon$. That is,
$\operatorname*{\mathbb{E}}_{\bar{\theta}}[\mathrm{ver}_{V,\mu}(\bar{\theta)}]\geq 1-\varepsilon.$
Proof.
Our strategy is to cast TL as Stochastic Gradient Ascent (SGD). We follow Shalev-Shwartz and Ben-David [2014, Section 14.3], which is presented for Descent (SGD) but is equivalent up to sign change.
Let $\varepsilon$ such that $B_{\varepsilon}$ and $\rho_{\varepsilon}$ are finite be given. Since $B_{\varepsilon}<\infty$, let $\theta^{\ast}$ be such that $\mathrm{ver}_{V,\mu}(\theta^{\ast})\geq 1-\varepsilon/2$ and $||\theta^{\ast}||\leq B_{\varepsilon}$. To prove the theorem, it suffices to prove that
$\operatorname*{\mathbb{E}}[\mathrm{ver}_{V,\mu}(\bar{\theta})]\geq\mathrm{ver}_{V,\mu}(\theta^{\ast})-\varepsilon/2.$
Following the notation in Algorithm 1, for every iteration $i\in[N]$, $r\in[R]\cup\{0\}$ and $s\in S(r)$, the definition of $\rho_{\varepsilon}$ gives $||\vec{d}_{s}(\theta_{i})||\leq\rho_{\varepsilon}$. Thus, for each $i\in[N]$, we can bound the norm of the update step by
$\left\lVert a_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}b_{r}(\theta_{i})a_{s}(\theta_{i})\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert\\ =\left|a_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}b_{r}(\theta_{i})a_{s}(\theta_{i})\right)\right|\cdot\left\lVert\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert\\ \leq(R+1)\cdot L_{a}+\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\left\lVert\vec{d}_{s}(\theta_{i})\right\rVert\leq(R+1)\cdot(L_{a}+\rho_{\varepsilon}).$
For the above, we used the fact that $\alpha_{s}(\theta_{i}),\beta_{r}(\theta_{i})\leq 1$, the definition of the answer length $L_{a}$ ($|S(r)|=L_{a}$), and the triangle inequality on |||||$\cdot$||| | ⋅ | |. Therefore, by [Shalev-Shwartz and Ben-David, 2014, Theorem 14.8] and Lemma A.4, TL (Algorithm 1) satisfies
$\operatorname*{\mathbb{E}}_{\bar{\theta}}\left[\mathrm{agree}_{\mathcal{T}^{\ast}_{V}}\left(\bar{\theta}\right)\right]\geq\mathrm{agree}(\theta^{\ast})-\varepsilon/2\geq 1-\varepsilon,$
where the right inequality is by the choice of $\theta^{\ast}$. The proof follows by observing that, for any $\bar{\theta}$ (and in particular an expected one), it holds that $\mathrm{agree}_{\mathcal{T}^{\ast}_{V}}(\bar{\theta})\leq\mathrm{ver}_{V,\mu}(\bar{\theta})$; this is because, for any $x$, whenever the transcript generated by $\mathcal{T}^{\theta}(x)$ agrees with $\pi^{\ast}$, then the verifier accepts (by definition of $\pi^{\ast}$).
A.2 Reinforcement Learning from Verifier Feedback
Our second learning method, Reinforcement Learning from Verifier Feedback (RLVF, Algorithm 2), does not require access to an honest transcript generator. Instead, the learner learns $P_{\theta}$ generates transcripts herself by emulating the interaction of the verifier with the current Self-Proving model $P_{\theta}$. When an accepting transcript is generated, the learner updates the parameters $\theta$ towards generating such transcript.
Hyperparameters: Learning rate $\lambda\in(0,1)$ and number of samples $N\in\mathbb{N}$.
Input: An autoregressive model family $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$, initial parameters $\theta_{0}\in\mathbb{R}^{d}$, verifier specification (code) $V$, and sample access to an input distribution $\mu$.
Output: A vector of parameters $\bar{\theta}\in\mathbb{R}^{d}$.
1
21:for $i=0,\dots,N-1$ do
2:Sample $x\sim\mu$.
3:Initialize $a_{0}\coloneqq y$ and $d_{i}\coloneqq\vec{0}$.
3      4:foreach $r=1,\dots R$ do
5:Sample the $r^{\text{th}}$ query # Emulate the verifier
$q_{r}\sim V_{q}(x,a_{0},q_{1},a_{1},\dots,q_{r},a_{r})$
6:Sample the $r^{\text{th}}$ answer # Forwards pass
$a_{r}\sim P_{\theta}(x,a_{0},q_{1},a_{1},\dots,q_{r},a_{r},q_{a_{r+1}})$
4            7:Let $\tau_{r}\coloneqq(a_{0},q_{1},\dots,a_{r-1},q_{r})$. for $s\in[L_{a}]$ do
5                  8:Let $a_{r,s}$ denote the $s^{\text{th}}$ token in $a_{r}$. Compute # Backwards pass
$\vec{d}_{s}(\theta_{i})$ $\coloneqq\nabla_{\theta}\log\Pr_{\sigma\sim p_{\theta_{i}}(x\tau_{r})}[\sigma=a_{r,s}].$
6      9:if $V(x,y,q_{1},a_{1},\dots,q_{R},a_{R})\ \mathrm{accepts}$ then
7            10:Update
$\theta_{i+1}\coloneqq\theta_{i}+\lambda\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\vec{d}_{s}(\theta_{i}).$
11:Output $\bar{\theta}\coloneqq\frac{1}{N}\sum_{i\in[N]}\theta_{i}$.
Algorithm 2 Reinforcement Learning from Verifier Feedback (RLVF)
Before we continue with formal analysis of Algorithm 2, let us make a few observations.
Firstly, the parameters are updated (line 11) only when an accepting transcript was generated. This means that the learner can first fully generate the transcript (lines 6-7), and then take backwards passes (line 9) only if the transcript was accepted by $V$. This is useful in practice (e.g. when using neural models) as backwards passes are more computationally expensive than forwards passes.
On the other hand, this means that RLVF requires the parameter initialization $\theta_{0}$ to have Verifiability bounded away from 0, so that accepting transcripts are sampled with sufficient probability. Fortunately, such a Self-Proving base model can be learned using TL. This gives a learning paradigm in which a somewhat-Self-Proving base model is learned with TL (with Verifiability $\delta>0$), and then “amplified” to a fully Self-Proving model using RLVF. This can be seen as an adaptation of the method of Nair et al. [2018] to the setting of Self-Proving models.
Secondly, in comparing Algorithms 1 and 2, we see that the latter (RLVF) does not keep track of the probabilities $\alpha_{s}$ and $\beta_{r}$. This is because, in RL terms, RLVF is an on-policy algorithm; it generates transcripts using the current learned model, unlike TL which samples them from a distribution whose parameterization is unknown to the learner. Hence, the update step in RLVF is simpler than TL. Furthermore, the RLVF learner does not require access to the density function of the query generator $V_{q}$ (Definition A.2) unlike its TL counterpart.
We now prove that the update step in RLVF maximizes the Verifiability of $P_{\theta}$; this is analogous to Lemma A.4 for TL. We leave it for future work to use Lemma A.6 to obtain convergence bounds on RLVF (analogous to Theorem A.5). As mentioned in Section 4.2, the gap between the lemma and a full convergence theorem (informally) reduces to the problem of obtaining convergence bounds for Policy Gradient methods, a challenging and active research direction (e.g. Agarwal et al. 2021).
Lemma A.6.
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. For any transcript $(x,y,q_{1},\dots,a_{R})$ we let $\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})$ denote the indicator random variable which equals 1 if and only if $V$ accepts the transcript. For any model $P_{\theta}$, denoting by $\mathrm{ver}_{V,\mu}(\theta)$ the verifiability of $P_{\theta}$ (Definition 3.3), it holds that
$\nabla_{\theta}\mathrm{ver}_{V,\mu}(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\\ (q_{r},a_{r})_{r=1}^{R}\end{subarray}}\left[\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\vec{d}_{s}(\theta)\right]$
where $(q_{r},a_{r})_{r=1}^{R}$ are as sampled in lines 5-6 of Algorithm 2, and $\vec{d}_{s}(\theta)$ is as defined in line 8 therein.
Proof.
Recall the transcript generator of $P^{\theta}$, denoted by $T^{\theta}_{V}$ (see Lemma A.4). By the definitions of Verifiability in Definition 3.3 and $V(x,y,q_{1},\dots,a_{R})$ in the lemma statement,
$\mathrm{ver}_{V,\mu}(\theta)$ $\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\end{subarray}}\left[V^{P_{\theta}}(x,y)\ \mathrm{accepts}\right]$
$=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\\ (q_{r},a_{r})_{r=1}^{R}\end{subarray}}\left[\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})\right]$
$=\operatorname*{\mathbb{E}}_{x\sim\mu}\left[\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}}[\mathrm{Acc}_{V}(x,\pi)\ ]\right]$ (11)
Now, for every input $x$, let $\Pi^{\ast}(x)\subset\Sigma^{\ast}$ denote the set of accepting transcripts:
$\Pi^{\ast}(x)\coloneqq\left\{\pi^{\ast}\in\Sigma^{\ast}:\mathrm{Acc}_{V}{x,\pi^{\ast}}\ \mathrm{accepts}\right\}.$
Noting that $\Pi^{\ast}(x)$ has finite or countably infinite cardinality, for any fixed input $x$ we can write
$\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}}[\mathrm{Acc}_{V}(x,\pi)]=\sum_{\pi^{\ast}\in\Pi^{\ast}(x)}\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}[\pi=\pi^{\ast}].$ (12)
We will use Equations (4) through (9) in the proof of Lemma A.4. Up to a change in index notation, these show that, for any $\pi^{\ast}$,
$\nabla_{\theta}\Pr_{\pi\sim\mathcal{T}^{\theta}(x)}[\pi=\pi^{\ast}]=\Pr_{\pi\sim\mathcal{T}^{\theta}(x)}[\pi=\pi^{\ast}]\cdot\sum_{\begin{subarray}{c}r\in R\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\nabla_{\theta}\vec{d}_{s}(\theta).$
Combining Equations 11 and 12, by linearity of expectation we have that
$\nabla_{\theta}\mathrm{ver}_{V,\mu}(\theta)$ $=\sum_{\pi^{\ast}\in\Pi^{\ast}(x)}\nabla_{\theta}\Pr_{\pi\sim\mathcal{T}^{\theta}(x)}[\pi=\pi^{\ast}]$
$=\operatorname*{\mathbb{E}}_{x\sim\mu}\left[\sum_{\pi^{\ast}\in\Pi^{\ast}(x)}\Pr_{\pi\sim\mathcal{T}^{\theta}(x)}[\pi=\pi^{\ast}]\cdot\sum_{\begin{subarray}{c}r\in R\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\nabla_{\theta}\vec{d}_{s}(\theta)\right]$
$=\operatorname*{\mathbb{E}}_{x\sim\mu}\left[\operatorname*{\mathbb{E}}_{\pi\sim\mathcal{T}^{\theta}(x)}\left[\mathrm{Acc}_{V}(x,\pi)\cdot\sum_{\begin{subarray}{c}r\in R\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\nabla_{\theta}\vec{d}_{s}(\theta)\right]\right]$
$=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi\sim\mathcal{T}^{\theta}(x)\end{subarray}}\left[\mathrm{Acc}_{V}(x,\pi)\cdot\sum_{\begin{subarray}{c}r\in R\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\nabla_{\theta}\vec{d}_{s}(\theta)\right]$
$=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\\ (q_{r},a_{r})_{r=1}^{R}\end{subarray}}\left[\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})\cdot\sum_{\begin{subarray}{c}r\in R\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\nabla_{\theta}\vec{d}_{s}(\theta)\right],$
where in the last equality, the probability is over $(q_{r},a_{r})$ sampled as in Algorithm 2, and it follows from the definition of the transcript generator $\mathcal{T}^{\theta}(x)$. ∎
Appendix B Annotations
We formally capture the modification described in Section 4.3 by introducing a transcript annotator and an answer extractor incorporated into the training and inference stages, respectively.
Fix a verifier $V$ in an $R$-round proof system with question length $L_{q}$ and answer length $L_{a}$. An annotation system with annotation length $\widetilde{L_{a}}$ consists of a transcript annotator $A$, and an answer extractor $E$.
In terms of efficiency, think of the annotator as an algorithm of the same computational resources as an honest prover in the system (see Definition 3.2, and the answer extractor as an extremely simple algorithm (e.g., trim a fixed amount of tokens from the annotation).
To use an annotation system the following changes need to be made:
  • At training time, an input $x$ and transcript $\pi$ is annotated to obtain $\widetilde{\pi}\coloneqq A(x,\pi)$, e.g. before the forwards backwards pass in TL (line 3 in Algorithm 1).
  • At inference time (i.e., during interaction between $V$ and $P_{\theta}$), the prover keeps track of the annotated transcript, but in each round passes the model-generated (annotated) answer through the extractor $E$ before it is sent to the verifier. That is, in each round $r\in[R]$, the prover samples
    $\widetilde{a_{r}}\sim P_{\theta}(x,y,q_{1},\widetilde{a_{1}},\dots,\widetilde{a_{r-1}},q_{r}).$
    The prover then extracts an answer $a_{r}\coloneqq E(\widetilde{a_{r}})$ which is sent to the verifier.
Appendix C A simple proof system for the GCD
The Euclidean algorithm for computing the Greatest Common Divisor (GCD) of two integers is possibly the oldest algorithm still in use today [Knuth, 1969]. Its extended variant gives a simple proof system.
Before we dive in, let us clarify what we mean by a proof system for the GCD. Paul has two integers $212$ and $159$; he claims that $GCD(212,159)=53$. An inefficient way for Veronica to check Paul’s answer is by executing the Euclidean algorithm on $(212,159)$ and confirm that the output is $53$. In an efficient proof system, Veronica asks Paul for a short string $\pi^{\ast}$ (describing two integers) with which she can easily compute the answer—without having to repeat Paul’s work all over. On the other hand, if Paul were to claim that “$GCD(212,159)=51$” (it does not), then for any alleged proof $\pi$, Veronica would detect an error and reject Paul’s claim.
The verifier in the proof system relies on the following fact.
Claim C.1 (Bézout’s identity [Bezout, 1779]).
Let $x_{0},x_{1}\in\mathbb{N}$ and $z_{0},z_{1}\in\mathbb{Z}$. If $z_{0}\cdot x_{0}+z_{1}\cdot x_{1}$ divides both $x_{0}$ and $x_{1}$, then $z_{0}\cdot x_{0}+z_{1}\cdot x_{1}=GCD(x_{0},x_{1})$.
Any coefficients $z_{0},z_{1}$ satisfying the assumption of C.1 are known as Bézout coefficients for $(x_{0},x_{1})$. C.1 immediately gives our simple proof system: For input $x=(x_{0},x_{1})$ and alleged GCD $y$, the honest prover sends (alleged) Bézout coefficients $(z_{0},z_{1})$. The Verifier accepts if and only if $y=z_{0}\cdot x_{0}+z_{1}\cdot x_{1}$ and $y$ divides both $x_{0}$ and $x_{1}$.
In this proof system the Verifier does not need to make any query; to fit within Definition 3.2, we can have the verifier issue a dummy query. Furthermore, by C.1 it is complete and has soundness error $s=0$. Lastly, we note that the Verifier only needs to perform two multiplications, an addition, and two modulus operations; in that sense, verification is more efficient than computing the GCD in the Euclidean algorithm.
Annotations.
To describe how a proof $z=(z_{0},z_{1})$ is annotated, let us first note how it can be computed. The Bézout coefficients can be found by an extension of the Euclidean algorithm. It is described in Algorithm 3.999Our description is the same as https://en.wikipedia.org/wiki/Extended_Euclidean_algorithm.
Input: Nonzero integers $x_{0},x_{1}\in\mathbb{N}$.
Output: Integers $(y,z_{0},z_{1})$, such that $y=GCD(x_{0},x_{1})$ and $(z_{0},z_{1})$ are Bézout coefficients for $(x_{0},x_{1})$.
1:Initialize $r_{0}=x_{0}$, $r_{1}=x_{1}$, $s_{0}=1$, $s_{1}=0$, and $q=0$.
12:while $r_{1}\neq 0$ do
3:Update $q\coloneqq(r_{0}//r_{1})$, where ///// / denotes integer division.
4:Update $(r_{0},r_{1})\coloneqq(r_{1},r_{0}-q\times r_{1})$.
2      5:Update $(s_{0},s_{1})\coloneqq(s_{1},s_{0}-q\times s_{1})$.
6:Output GCD $y=r_{0}$ and Bézout coefficients $z_{0}\coloneqq s_{0}$ and $z_{1}\coloneqq(r_{0}-s_{0}\cdot x_{0})/x_{1}$.
Algorithm 3 Extended Euclidean algorithm
Referring to Algorithm 3, the annotation of a proof $z=(z_{0},z_{1})$ will consist of intermediate steps in its computation. Suppose that in each iteration of the While-loop, the algorithm stores each of $r_{0}$, $s_{0}$ and $q$ in an arrays $\vec{r_{0}}$, $\vec{s_{0}}$ and $\vec{q}$. The annotation $\tilde{z}$ of $z$ is obtained by concatenating each of these arrays. In practice, to avoid the transformer block (context) size from growing too large, we fix a cutoff $T$ and first trim each array to its first $T$ elements.
We formalize this in the terminology of Appendix B by defining a Transcript Annotator and Answer Extractor. Note that, since our proof system consists only of one “answer” $z$ send from the prover to the verifier, the entire transcript $\pi$ is simply $z=(z_{0},z_{1})$. [^10] Such a proof system is known as an NP-proof system and defines the complexity class NP [Karp, 1972].
  • Transcript Annotator $A$: For a fixed cutoff $T$ and given input $x=(x_{0},x_{1})$ and transcript $z=(z_{0},z_{1})$, $A$ executes Algorithm 3 on input $x=(x_{0},x_{1})$. During the execution, $A$ stores the first $T$ intermediate values of $r_{0}$, $s_{0}$ and $q$ in arrays $\vec{r_{0}}$, $\vec{s_{0}}$ and $\vec{q}$. It outputs $A(x,z)\coloneqq(\vec{r_{0}},\vec{s_{0}},\vec{q},z)$.
  • Answer Extractor $E$: Given an annotated transcript $\tilde{z}=(\vec{r_{0}},\vec{s_{0}},\vec{q},z)$, outputs $E(\tilde{z})\coloneqq z$.
We note that the computational complexity of $A$ is roughly that of the honest prover, i.e., Algorithm 3 (up to additional space due to storing intermediate values). As for $E$, it can be implemented in logarithmic space and linear running time in $|\tilde{z}|$, i.e., the length of the description. [^11] That is, if integers are represented by $n$-bits, then $E$ has space complexity $O(\log n+\log T)$ and running time $O(n\cdot T)$.
Appendix D Experiment details
We provide details of how we implemented the experiments in Section 5 and additional figures for each experiment.
Model architecture.
We use Karpathy’s nanoGPT. [^12] https://github.com/karpathy/nanoGPT We use a 6.7M parameter architecture of $8$ layers, $8$ attention heads, and $256$ embedding dimensions. We optimized hyperparameters via a random hyperparameter search, arriving at learning rate $0.0007$, AdamW $\beta_{1}=0.733$ and $\beta 2=0.95$, $10\%$ learning rate decay factor, no dropout, gradient clipping at $2.0$, no warmup iterations, and $10\%$ weight decay.
Data.
We sample integers from the $\log_{10}$-uniform distribution over $\{1,\dots,10^{4}\}$. Models in Tables 1 and 2 are trained for 100K iterations on a dataset of of $\approx$10M samples. For Figure 3 (base ablation) we train for 20K iterations on a dataset of $\approx$1M samples; this is because this setting required 68 many runs in total, whereas the annotation-cutoff ablation required 18 longer runs.
Compute.
All experiments were run on a machine with an NVIDIA A10G GPU, 64GB of RAM, and 32 CPU cores. Longer runs (annotation-cutoff ablation) took about 75 minutes each. Shorter runs (base ablation) took about 15 minutes. The total running time of our experiments was approximately 40 hours, excluding time dedicated to a random hyperparameter search. The overall disk space needed for our models and data (to be made available upon publication) is 4GB.
Representing integers.
We fully describe how integer sequences are encoded. As a running example, we will use base $210$. To encode a sequence of integers, each integer is encoded in base $210$, a sign is prepended and a delimiter is appended, with a unique delimiter identifying each component of the sequence. For example, consider the input integers $x_{0}=212$ (which is $12$ in base 210) and $x_{1}=159$. Their GCD is $y=53$, with Bézout coefficients $z_{0}=1$ and $z_{1}=-1$. Therefore, the sequence $(212,159,53,1,-1)$ is encoded as
+,1,2,x0,+,159,x1,+,53,y,+,1,z0,-,1,z1
where commas are added to distinguish between different tokens. Null tokens are appended to pad all sequences in a dataset to the same length. Both the input and the padding components are ignored when computing the loss and updating parameters.
Annotations
Annotations are encoded as above, with each component in an intermediate step $\pi_{t}$ delimited by a unique token. Since different integer pairs may require a different number of intermediate steps to compute the Bézout coefficients, we chose to pad all annotaitons to the same length $T$ by the last step $\pi_{T}$ in the sequence (which consists of the final Bézout coefficients). This ensures that the final component output by the model in each sequence should be the Bézout coefficient, and allows us to batch model testing (generation and evaluation) resulting in a 1000x speed-up over sequential testing.
As an example, consider the inputs $x_{0}=46$ and $x_{1}=39$. Tracing through the execution of Algorithm 3, we have
$\begin{array}[]{c|c|c|c|c|c|c|c|c}&x_{0}&x_{1}&y&\vec{s_{0}}&\vec{r_{0}}&\vec{q}&z_{0}&z_{1}\\ \hline\cr&46&39&&1&46&1&&\\ &&&&0&39&5&&\\ &&&&1&7&1&&\\ &&&&-5&4&1&&\\ &&&&6&3&3&&\\ &&&1&&&&-11&13\\ \end{array}$
To encode this as an annotated transcript for the transformer, we must specify a base of representation and an annotation cutoff. Suppose that we wish to encode this instance in base $B=10$ and cutoff $T=3$. Then the input with the annotated transcript is encoded as
+,4,6,x0,+,3,9,x1,+,1,y,
+,1,z0’,+,4,6,z1’,+,1,q’,
+,0,z0’’,+,3,9,z1’’,+,5,q’’
+,1,z0’’’,+,7,z1’’’,+,1,q’’’,
-,1,1,z0,+,1,3,z1
where commas are used to separate between tokens, and linebreaks are added only for clarity. Notice the three types of tokens: signs, digits, and delimiters. Notice also that the output $y$ is added immediately after the input, followed by the annotated transcript (whose six tokens comprise the proof itself). Since the Self-Proving model we train has causal attention masking, placing the output $y$ before the proof means that the model “commits” to an output and only then proves it.
Appendix E Additional figures for Section 5
Figure 4: Verifiability as a function of the number of samples $N$. Each iteration (X axis) is a batch of 1024 samples from a dataset of $\approx$10M sequences. Every 10k iterations, Verifiability was evaluated on a held-out dataset of 1k inputs (as described in Section 5). $T$ is the number of steps in Annotated Transcript Learning (Figure 2), and $T=0$ is non-annotated Transcript Learning. Each $T$ was run with three seeds, with mean depicted by the curve and standard error by the shaded area.
Figure 5: Full version of Figure 2, with all $T\in\{0,3,4,5,6,7\}$ possibilities for the annotation cutoff in Annotated Transcript Learning ($T=0$ means no annotation). For a zoom-in of the right end of the axis, see Figure 2.